OpenStudy (anonymous):

Which of the following inequalities matches the graph? graph of an inequality with a dashed line through the points (0, 7) and (1, 5) with shading below the line 2x + y < 7 2x - y > 7 2x + y < -7 The correct inequality is not listed

OpenStudy (anonymous):

first, start with the slope: 7-5=2 0-1=-1 2/-1=-2

OpenStudy (anonymous):

plug in a point to find b: 7=0(-2)+b 7=b

OpenStudy (anonymous):

now write an equation: -2x+7<y

OpenStudy (anonymous):

now add 2x to each side 7<2x+y